solve each equation.
5 + 4(t - 2) = 2(t + 7) + 1

Answer by TimothyLamb(4379)

5 + 4(t - 2) = 2(t + 7) + 1
5 + 4t - 8 = 2t + 14 + 1
4t - 2t = 14 + 1 + 8 - 5
2t = 18
t = 9
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php